Find the length of each side of a triangle if the perimeter is 73. The length of the second side is twice the length of the first side and the length of the third side is five more than the length of the first side. I have no idea!! : Find the length of each side of a triangle if the perimeter is 73. The length of the second side is twice the length of the first side and the length of the third side is five more than the length of the first side. I have no idea!!
Answer by dddiwakar_2004(2)

solution-
Given the Perimeter of a triangle is 73
Let the first side be x
second side = 2x
third side = x+5
PERIMETER OF A TRIANGLE = SUM OF ALL SIDES
73=X+2X+X+5
73=4X+5
73-5=4X+5-5
68=4X
68/4=4X/4
17=X
THIS IMPLISE THAT FIRST SIDE IS X = 17
SECOND SIDE IS 2X= 2*17=34
THIRD SIDE IS X+5=17+5=22
FOR MORE HELP VISIT : WWW.TAKEATUITION.COM